A gas occupies a volume of 500 mL at 300K. To what temperature should the gas be lowered to bring the volume to 300 mL?


1
Expert's answer
2021-01-25T04:24:32-0500

"V_1 = 500 \;mL \\

T_1 = 300 \;K \\

V_2 = 300 \;mL \\

P = constant \\

\frac{V_1}{T_1} = \frac{V_2}{T_2} \\

T_2 = \frac{V_2 \times T_1}{V_1} \\

= \frac{300 \times 300}{500} \\

= 180 \;K"

Answer: 180 K